OpenStudy (anonymous):

Solve the following equation for x: logx + log(x+2) = log(3x) a. x = 0 b. x = 1 c. x = 0,1 d. There are no solutions.

OpenStudy (anonymous):

logx + log(x+2) = log(3x) log(x(x+2))= log(3x) log(x^2+2x)= log(3x) x^2+2x=3x x^2-x=0 x(1-x)=0 x = 0,1